OpenStudy (anonymous):

X^2-3x-10/x+2 A. X-5; where x=-2 B. X-5; where x=5 C. 1/x-2 D. X-2

OpenStudy (muzzack):

Step by step solution : Step 1 : 10 Simplify x2-3x - —— x Rewriting the whole as an Equivalent Fraction : 1.1 Subtracting a fraction from a whole Rewrite the whole as a fraction using x as the denominator : x2 - 3x (x2 - 3x) • x x2 - 3x = ——————— = ————————————— 1 x Equivalent fraction : The fraction thus generated looks different but has the same value as the whole Common denominator : The equivalent fraction and the other fraction involved in the calculation share the same denominator Pulling out like terms : 1.2 Pull out like factors : x2 - 3x = x • (x - 3) Adding fractions that have a common denominator : 1.3 Adding up the two equivalent fractions Add the two equivalent fractions which now have a common denominator Combine the numerators together, put the sum or difference over the common denominator then reduce to lowest terms if possible: x • (x-3) • x - (10) x3 - 3x2 - 10 ———————————————————— = ————————————— x x Equation at the end of step 1 : (x3 - 3x2 - 10) ——————————————— + 2 x Step 2 : x3-3x2-10 Simplify ————————— + 2 x Rewriting the whole as an Equivalent Fraction : 2.1 Adding a whole to a fraction Rewrite the whole as a fraction using x as the denominator : 2 2 • x 2 = — = ————— 1 x Polynomial Roots Calculator : 2.2 Find roots (zeroes) of F(x) = x3 - 3x2 - 10 Polynomial Roots Calculator is a set of methods aimed at finding values of x for which F(x)=0 Rational Roots Test is one of the above mentioned tools. It would only find Rational Roots that is numbers x which can be expressed as the quotient of two integers The Rational Root Theorem states that if a polynomial zeroes for a rational number P/Q then P is a factor of the Trailing Constant and Q is a factor of the Leading Coefficient In this case, the Leading Coefficient is 1 and the Trailing Constant is -10. The factor(s) are: of the Leading Coefficient : 1 of the Trailing Constant : 1 ,2 ,5 ,10 Let us test .... P Q P/Q F(P/Q) Divisor -1 1 -1.00 -14.00 -2 1 -2.00 -30.00 -5 1 -5.00 -210.00 -10 1 -10.00 -1310.00 1 1 1.00 -12.00 2 1 2.00 -14.00 5 1 5.00 40.00 10 1 10.00 690.00 Polynomial Roots Calculator found no rational roots Adding fractions that have a common denominator : 2.3 Adding up the two equivalent fractions (x3-3x2-10) + 2 • x x3 - 3x2 + 2x - 10 ——————————————————— = —————————————————— x x Checking for a perfect cube : 2.4 x3 - 3x2 + 2x - 10 is not a perfect cube Trying to factor by pulling out : 2.5 Factoring: x3 - 3x2 + 2x - 10 Thoughtfully split the expression at hand into groups, each group having two terms : Group 1: 2x - 10 Group 2: x3 - 3x2 Pull out from each group separately : Group 1: (x - 5) • (2) Group 2: (x - 3) • (x2) Bad news !! Factoring by pulling out fails : The groups have no common factor and can not be added up to form a multiplication. Polynomial Roots Calculator : 2.6 Find roots (zeroes) of F(x) = x3 - 3x2 + 2x - 10 See theory in step 2.2 In this case, the Leading Coefficient is 1 and the Trailing Constant is -10. The factor(s) are: of the Leading Coefficient : 1 of the Trailing Constant : 1 ,2 ,5 ,10 Let us test .... P Q P/Q F(P/Q) Divisor -1 1 -1.00 -16.00 -2 1 -2.00 -34.00 -5 1 -5.00 -220.00 -10 1 -10.00 -1330.00 1 1 1.00 -10.00 2 1 2.00 -10.00 5 1 5.00 50.00 10 1 10.00 710.00 Polynomial Roots Calculator found no rational roots Final result : x3 - 3x2 + 2x - 10 —————————————————— x